# Uncertainty-Informed Active Perception for Open Vocabulary Object Goal Navigation

### v1

@@ Abstract
Mobile robots exploring indoor environments increasingly rely on vision-language models to perceive high-level semantic cues in camera images, such as object categories. Such models offer the potential to substantially advance robot behaviour for tasks such as object-goal navigation (ObjectNav), where the robot must locate objects specified in natural language by exploring the environment. Current ObjectNav methods heavily depend on prompt engineering for perception and do not address the semantic uncertainty induced by variations in prompt phrasing. Ignoring semantic uncertainty can lead to suboptimal exploration, which in turn limits performance. Hence, we propose a semantic uncertainty-informed active perception pipeline for ObjectNav in indoor environments. We introduce a novel probabilistic sensor model for quantifying semantic uncertainty in vision-language models and incorporate it into a probabilistic geometric-semantic map to enhance spatial understanding. Based on this map, we develop a frontier exploration planner with an uncertainty-informed multi-armed bandit objective to guide efficient object search. Experimental results demonstrate that our method achieves ObjectNav success rates comparable to those of state-of-the-art approaches, without requiring extensive prompt engineering.
@@ I Introduction
Robots benefit from goal-directed exploration in human-centric environments to accomplish tasks such as locating and delivering objects, as well as assisting with household activities. These tasks require a robot to detect objects and understand spatial layouts with task-relevant contextual semantic cues. To this end, grounding language instructions in perceptual and spatial representations of the scene, also known as open-vocabulary perception, is key to efficient exploration and success of the task. However, open-vocabulary perception is inherently uncertain due to the wide variety of objects that exist in households and the many ways they can be described in natural language. Quantifying this perception uncertainty and accounting for it during exploration is an important element to realise reliable task execution.
@@
A common strategy employed by ObjectNav methods is to construct a 2D map of the environment online using observed geometric information, such as spatial layout and obstacles. Geometric maps enable tracking explored areas and guiding exploration toward map corners [15] or frontiers between known and unknown space, known as frontier-based exploration [26]. Semantic information, such as object categories, can also be integrated into maps using deep learning-based semantic segmentation [4], which enables exploration of areas with semantic similarity to the target object. However, most approaches are closed-vocabulary, i.e. limited to a fixed set of objects defined at training time and do not scale well to the wide variety of objects found in the real world. State-of-the-art approaches for ObjectNav [28, 5, 6, 16] leverage open-vocabulary perception using vision-language models (VLMs) [18, 12, 11], allowing for semantically informed exploration scalable to potentially infinite object categories. Trained on large-scale image-text pairs, VLMs provide semantic relevance between arbitrary user-specified text prompts and images. However, VLMs are often sensitive to prompt phrasing. Variations in prompts, such as “chair,” “a chair,” or “there is a chair nearby” produce different semantic relevance scores from VLMs, given the same image. This can lead to inconsistency in navigating to the target object in the ObjectNav scenario. Existing approaches [5, 28] often rely on fixed hand-tuned prompts appended to object names, which ignore this inconsistency and do not address this semantic uncertainty in VLM predictions. Ignoring semantic uncertainties is suboptimal and does not exploit the full potential of planning algorithms. Our paper aims to overcome this within the ObjectNav problem.
@@ II Related Work
In contrast, modular ObjectNav methods [6, 28, 32, 29, 5, 31] decompose the task into perception, mapping, exploration and point-goal navigation. These approaches build a map of the environment using geometric and semantic information. A planner selects waypoints for exploration of the environment, while point-goal navigation determines the actions required to reach them.
@@ IV Our Approach
Our uncertainty-informed ObjectNav approach explores the environment based on probabilistic estimates of semantic relevance. Our approach enables a robot to navigate to user-specified objects in an unknown 3D environment based on open-vocabulary perception, conceptually depicted in Fig. 2. We first perform object detection in each received camera frame as described in Sec. IV-A. If the target object is detected in the frame, the robot directly navigates to it. Otherwise, we employ an active exploration strategy that identifies semantically relevant regions, such as a “kitchen” for target objects like “cup”, by exploiting semantic correspondences and accounting for uncertainty using our probabilistic sensor model described in Sec. IV-B. We construct a geometric-semantic grid map based on the new sensor model, updating occupancy probabilities and semantic relevance information online, as discussed in Sec. IV-C. Our multi-arm bandit frontier-based planner uses our geometric-semantic map to guide the robot to regions with high semantic relevance, as outlined in Sec. IV-D.
@@ IV-A Object Detection and Point Goal Navigation
We follow the approach of Yokohama et al. [28] for object detection. Each recorded RGB image is processed by the object detection module, which attempts to detect the target object $o\in\mathcal{O}$ in the frame using a combination of two object detectors. We rely on YOLOv7 [22], a supervised learning-based detector trained a pre-defined closed set of $91$ object categories $\mathcal{O}$, including a wide range of outdoor and indoor scenes. To allow our approach to work with potentially infinite and a priori unknown user-defined object categories of interest, we further use Grounding DINO [14] as an open-vocabulary object detector capable of detecting a wide range of objects specified by arbitrary text prompts. If the target object $o$ is detected, we use MobileSAM [30] to segment $o$ in the RGB image. To estimate the goal pose $\mathbf{x}^{w}_{g}$, the segmentation mask is applied to the depth image and projected into 3D world coordinates using known camera parameters. We use the centroid of the projected depth information of $o$ to determine a target object position $\mathbf{x}^{w}_{g}\in\mathcal{G}$. Then, we use DD-PPO [24] to navigate to $\mathbf{x}^{w}_{g}$ from the current pose of the robot $\mathbf{x}^{w}_{t}$.
@@ V-A Experimental Setup
Datasets. To assess the effectiveness of our ObjectNav approach, we conduct experiments in realistic indoor environments using the Habitat simulator [20], a popular simulator for ObjectNav evaluations [2]. We use two datasets, Matterport3D (MP3D) [3] and Habitat-Matterport 3D (HM3D) [19], both offering complex 3D reconstructions of indoor spaces. MP3D provides 3D scans with rich semantic annotations but limited unique environments and data imperfections. HM3D offers 1,000 high-quality scans with complex layouts and varied conditions. We evaluate our approach across varied, realistic scenes in 2,195 MP3D and 2,000 HM3D episodes, which contain 20 scenes, six object categories and 11 scenes and 21 object categories, respectively, following prior work [28].
@@ V-C Prompt Ensembling for Semantic Relevance
V-D Performance on the ObjectNav Task
Our results are summarized in Tab. II. Frontier-based Closest-FBE planning performs worst, often trapping the robot in narrow, nearby regions. Frontier-based Random-FBE avoids this by selecting random frontiers, enabling more thorough exploration of unknown space. However, neither approach is semantically informed but purely geometric, resulting in suboptimal performance. Among the semantically-informed methods, in line with prior works, VLFM outperforms CoW, ESC, and ZSON, due to the strength of its underlying VLM and updating the map with cosine similarity values for each received camera frame. Our uncertainty-informed planners achieve success rates comparable to VLFM, despite not relying on hand-crafted prompts. Particularly, I-FBE2 outperforms VLFM across many prompt variations on HM3D as shown in Tab. I. We demonstrate that open-vocabulary ObjectNav methods relying on a single hand-tuned prompt are brittle and show that our uncertainty-informed planners leveraging our probabilistic sensor model offer a more reliable and robust alternative. However, our methods perform slightly worse on the SPL metric than VLFM, which might be due to additional exploratory steps taken under high semantic relevance uncertainty.
@@ V-D Performance on the ObjectNav Task
VI Conclusion
In this work, we proposed a training-free, open-vocabulary, semantic uncertainty-informed active perception pipeline for the ObjectNav problem in mobile robotics. We introduce a novel probabilistic sensor model based on prompt ensembles to estimate semantic relevance and uncertainty from VLMs. Our framework incorporates a probabilistic geometric-semantic map and uncertainty-informed frontier planners to tackle the brittleness of existing ObjectNav approaches relying on a single fixed hand-engineered prompt. Our experimental results show that the success of prior open-vocabulary ObjectNav relies on extensive prompt engineering. We further show that semantic relevance scores from a VLM prompt-ensemble are approximately normally distributed and validate the design of our probabilistic sensor model. Our uncertainty-informed planner performs comparably to state-of-the-art open-vocabulary ObjectNav approaches that rely on fixed hand-engineered prompts. Overall, our findings demonstrate how VLM uncertainty affects open-vocabulary ObjectNav and that our approach enables competitive ObjectNav performance while reducing dependence on prompt engineering.

### v2

@@ Abstract
Mobile robots exploring indoor environments increasingly rely on vision-language models to perceive high-level semantic cues in camera images, such as object categories. Such models offer the potential to advance robot behaviour for tasks such as object-goal navigation (ObjectNav), where the robot must locate objects specified in natural language by exploring the environment. Current ObjectNav methods focus on prompt engineering for perception and do not address the semantic uncertainty induced by variations in prompt phrasing. Ignoring semantic uncertainty can lead to suboptimal exploration, which in turn limits performance. Hence, we propose a semantic uncertainty-informed active perception pipeline for ObjectNav in indoor environments. We introduce a novel probabilistic sensor model for quantifying semantic uncertainty in vision-language models and incorporate it into a probabilistic geometric-semantic map to enhance spatial understanding. Based on this map, we develop a frontier exploration planner with an uncertainty-informed multi-armed bandit objective to guide efficient object search. Experimental results demonstrate that our method achieves ObjectNav success rates comparable to those of state-of-the-art approaches, without requiring extensive prompt engineering.
@@ I Introduction
Robots benefit from effective goal-directed exploration in human-centric environments to accomplish tasks such as locating and delivering objects, as well as assisting with household activities. These tasks require a robot to detect objects and understand spatial layouts with task-relevant contextual semantic cues. To this end, grounding language instructions in perceptual and spatial representations of the scene, also known as open-vocabulary perception, is key to efficient exploration and success of the task. However, open-vocabulary perception is inherently uncertain due to the wide variety of objects that exist in households and the many ways they can be described in natural language. Quantifying this perception uncertainty and accounting for it during exploration is an important element to realise reliable task execution.
@@
A common strategy employed by ObjectNav methods is to construct a 2D map of the environment online using observed geometric information, such as spatial layout and obstacles. Geometric maps enable tracking explored areas and guiding exploration toward map corners [15] or frontiers between known and unknown space, known as frontier-based exploration [27]. Semantic information, such as object categories, can also be integrated into maps using deep learning-based semantic segmentation [4], which enables exploration of areas with semantic similarity to the target object. However, most approaches are closed-vocabulary, i.e. limited to a fixed set of objects defined at training time and do not scale well to the wide variety of objects found in the real world. State-of-the-art approaches for ObjectNav [29, 5, 7, 16] leverage open-vocabulary perception using vision-language models (VLMs) [18, 12], allowing for semantically informed exploration scalable to potentially infinite object categories. Trained on large-scale image-text pairs, VLMs provide semantic relevance between arbitrary user-specified text prompts and images. However, VLMs are often sensitive to prompt phrasing. Variations in prompts, such as “chair,” “a chair,” or “there is a chair nearby” produce different semantic relevance scores from VLMs, given the same image. This can lead to inconsistency in navigating to the target object in the ObjectNav scenario. Existing approaches [5, 29] often rely on fixed hand-tuned prompts appended to object names, which ignore this inconsistency and do not address this semantic uncertainty in VLM predictions. Ignoring semantic uncertainties is suboptimal and does not exploit the full potential of planning algorithms. Our paper aims to overcome this within the ObjectNav problem.
@@ II Related Work
In contrast, modular ObjectNav methods [7, 29, 33, 30, 5, 32] decompose the task into perception, mapping, exploration and point-goal navigation. Approaches build maps of the environment using geometric [22] and semantic information [6]. A planner selects waypoints for exploration of the environment, while point-goal navigation determines the actions required to reach them.
@@ IV Our Approach
Our uncertainty-informed ObjectNav approach explores the environment based on probabilistic estimates of semantic relevance. Our approach enables a robot to navigate to user-specified objects in an unknown 3D environment based on open-vocabulary perception, conceptually depicted in Fig. 2. We first perform object detection in each camera frame as described in Sec. IV-A. If the target object is detected, the robot directly navigates to it. Otherwise, we employ an active exploration strategy that identifies semantically relevant regions, such as a “kitchen” for target objects like “cup”, by exploiting semantic correspondences and accounting for uncertainty using our probabilistic sensor model described in Sec. IV-B. We construct a geometric-semantic grid map based on the new sensor model, updating occupancy and semantic relevance information online, as discussed in Sec. IV-C. Our multi-arm bandit frontier-based planner uses our geometric-semantic map to guide the robot to regions with high semantic relevance, as outlined in Sec. IV-D.
@@ IV-A Object Detection and Point Goal Navigation
We follow the approach of Yokohama et al. [29] for object detection. Each recorded RGB image is processed by the object detection module, which attempts to detect the target object $o\in\mathcal{O}$ in the frame using two object detectors. We rely on YOLOv7 [23], a supervised learning-based detector trained on a pre-defined closed set of $91$ object categories $\mathcal{O}$, including a wide range of outdoor and indoor scenes. To allow our approach to work with potentially infinite and a priori unknown user-defined object categories of interest, we further use Grounding DINO [14] as an open-vocabulary object detector capable of detecting a wide range of objects specified by arbitrary text prompts. If the target object $o$ is detected, we use MobileSAM [31] to segment $o$ in the RGB image. To estimate the goal pose $\mathbf{x}^{w}_{g}$, the segmentation mask is applied to the depth image and projected into 3D world coordinates using known camera parameters. We use the centroid of the projected depth information of $o$ to determine a target object position $\mathbf{x}^{w}_{g}\in\mathcal{G}$. Then, we use DD-PPO [25] to navigate to $\mathbf{x}^{w}_{g}$ from the current pose of the robot $\mathbf{x}^{w}_{t}$.
@@ V-A Experimental Setup
Datasets and hardware. To assess the effectiveness of our ObjectNav approach, we conduct experiments in realistic indoor environments using the Habitat simulator [20], a popular simulator for ObjectNav evaluations [2]. We use two datasets, Matterport3D (MP3D) [3] and Habitat-Matterport 3D (HM3D) [19], both offering complex 3D reconstructions of indoor spaces. MP3D provides 3D scans with rich semantic annotations but limited unique environments and data imperfections. HM3D offers 1,000 high-quality scans with complex layouts and varied conditions. We evaluate our approach across varied, realistic scenes in 2,195 MP3D and 2,000 HM3D episodes, which contain 20 scenes, six object categories and 11 scenes and 21 object categories, respectively, following prior work [29]. Experiments are conducted on a workstation with a 12th-generation Intel Core i7 CPU, 64 GB RAM, and an NVIDIA RTX A5000 GPU.
@@ V-C Prompt Ensembling for Semantic Relevance
V-D ObjectNav Performance
Our results are summarized in Tab. II. Frontier-based Closest-FBE planning performs worst, often trapping the robot in narrow, nearby regions. Frontier-based Random-FBE avoids this by selecting random frontiers, enabling more thorough exploration of unknown space. However, neither approach is semantically informed but purely geometric, resulting in suboptimal performance. Among the semantically-informed methods, in line with prior works, VLFM outperforms CoW, ESC, and ZSON, due to the strength of its underlying VLM and updating the map with cosine similarity values for each received camera frame. Our uncertainty-informed planners achieve success rates comparable to VLFM, despite not relying on hand-crafted prompts. Particularly, I-FBE2 outperforms VLFM across many prompt variations on HM3D as shown in Tab. I. We demonstrate that open-vocabulary ObjectNav methods relying on a single hand-tuned prompt are brittle. In contrast, our uncertainty-informed planners leveraging our probabilistic sensor model offer a more reliable and robust alternative. However, our method consistently underperforms slightly on the SPL metric compared to VLFM. This is primarily because our planner is designed to actively explore regions with high semantic relevance uncertainty. As a result, the agent occasionally replans toward uncertain regions even if they are not highly semantically relevant, leading to detours en route to the target object. On our evaluation setup, each replanning step takes approximately 420 ms of wall-clock time per timestep.
VI Conclusions and Future Work
In this work, we proposed a training-free, open-vocabulary, semantic uncertainty-informed active perception pipeline for the ObjectNav problem in mobile robotics. We introduce a novel probabilistic sensor model based on prompt ensembles to estimate semantic relevance and uncertainty from VLMs. Our framework incorporates a probabilistic geometric-semantic map and uncertainty-informed frontier planners to address the brittleness of existing ObjectNav approaches, which rely on a single, fixed, hand-engineered prompt. Our experimental results show that the success of prior open-vocabulary ObjectNav relies on extensive prompt engineering. In contrast, our uncertainty-informed planner performs comparably to state-of-the-art open-vocabulary ObjectNav approaches that rely on fixed hand-engineered prompts while reducing dependence on prompt engineering. Future work will investigate the real-world deployment of our approach in larger environments under sensor noise.